OpenStudy (anonymous):

i have the answer but the problem but no explanation.. i need to know how to solve this. how do i get there. thanks. 4x2 = -8x -1 the answer comes out to this fraction -2 ±√3 ________ 2

OpenStudy (anonymous):

and the problem... here is a picture just in case

OpenStudy (anonymous):

use the quadratic formula after getting all terms on one side 4x2+8x+1 so a=4, b=8, and c=1 plug into qf

OpenStudy (anonymous):

did you take a look at the picture? it didnt type out correctly. sorry. i really appreciate your help.

OpenStudy (anonymous):

I understood the problem..do you know the quadratic formula?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

ok I will type it out

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

\[-b \pm \sqrt{b ^{2}-4ac}/2a\]

OpenStudy (anonymous):

2a is in the denominator so substitute a, b, and c

OpenStudy (anonymous):

\[-8 \pm \sqrt{64-16}/ 8\]

OpenStudy (anonymous):

\[-8\pm \sqrt{48}/8 = -8\pm \sqrt{16*3} /8 -8\pm4\sqrt{3} /8\]

OpenStudy (anonymous):

I skipped the initial step of plugging in the values for a, b, and c do you get that part?

OpenStudy (anonymous):

yes i do thank you

OpenStudy (anonymous):

so do you understand it

OpenStudy (anonymous):

its for friend, but yes i think so. im passing the info along. thank you for your time.

OpenStudy (anonymous):

ok you're very welcome

OpenStudy (anonymous):

one last question.. how did you know that the qf was necessary? just in case other problems like these come up. how can i decipher whether i need it or not.

OpenStudy (anonymous):

if you can't solve the quadratic equation by factoring then your other options are by completing the square or the quadratic formula

OpenStudy (anonymous):

not sure if youre still there, but one thing my friend didnt understand was how it was simplified to the final answer. isnt the 4 outside the square root technically tied to it and you can't really simplify it?

OpenStudy (radar):

The simplification was valid. The 48 inside the radical was factored to 16 and 3, the square root of 16 was accomplished by moving the 4 outside of the radical leaving only the square root of 3 within the radical.

OpenStudy (radar):

The simplification of\[-8\pm(4\sqrt{3})/8=-4(2\pm \sqrt{3})/8=(-2\pm \sqrt{3})/2\]